Nicoley269:

how do i solve sqrt{a-2}=6

Nicoley269:

\[\sqrt{a-2}=6\]

Hero:

Step 1: Square both sides

Nicoley269:

a-2=36 a=38?

Hero:

To check, insert the value you found for \(a\) back into the original equation

Nicoley269:

its correct